Monthly food expenditures for families in Rodriguez Rizal averaged Php6,500 w


a standard deviation of Php750. Assuming that the monthly food expenditures are


normally distributed. What percentages of these expenditures are between Pho5.900


and Php7,8502

1
Expert's answer
2022-03-04T04:27:07-0500

In this case, we need to find the area between X = 5,900 and X =7,850


By using the Z score formula given by:


Z = "\frac{X -\mu}{\sigma}" 

Given "\mu" = 6,500 and "\sigma" = 750

When X = 5,900


Z = "\frac{5,900 -6,500}{750}"  = -0.8


When X = 7,850


Z = "\frac{7,850 -6,500}{750}"  = 1.8


Therefore, we need to find the area between Z = -0.8 and Z = 1.8

That is;

P(-0.8<Z<1.8) = P(Z<1.8) - P(Z<-0.8)

From the standard normal table,

P(Z<1.8) = 0.9641

Also,

P(Z<-0.8) = 0.2119

Therefore,

P(-0.8<Z<1.8) = P(Z<1.8) - P(Z<-0.8) = (0.9641 - 0.2119) = 0.7522


Answer: 0.7522